OpenStudy (anonymous):

What is the volume of a square prism whose base edges are 2 centimeters and height is 8 centimeters?

OpenStudy (anonymous):

Just find the base and multiply by the height in your case it is the same as finding the volume of a cube. \[Base \times Hieght = (2 \times 2 ) \times 8 = 32\]

OpenStudy (anonymous):

got it thanx